OpenStudy (anonymous):

Solve the equation. (Sort your answers in increasing order, first by their imaginary parts, then by their real parts.) x^4 + 81 = 0

OpenStudy (freckles):

you could factor x^4+81 \[a^2+b^2=(a+bi)(a-bi)\]

OpenStudy (freckles):

but it look like you have been asking trig questions so i have a better way

OpenStudy (anonymous):

this might seem like a stupid question but that is only for squared... what abt the rest of the problem?

OpenStudy (anonymous):

or would it be twice?

OpenStudy (freckles):

\[x^4=-81 \\ x^4=-81(\cos(0)+i \sin(0)) \\ \text{ so since cos and sin are periodic we have } \\x^4=-81(\cos(0+2n \pi)+i \sin(0+2n \pi))\] you want 4 answers so find your first answer when n=0 next answer when n=1 next n=2 next n=3 that is 4 roots

OpenStudy (anonymous):

\[x^4+9^2+18x^2-18x^2=0\] \[\left( x^2+9 \right)^2-\left( 3\sqrt{2}x \right)^2=0\] \[\left( x^2+3\sqrt{2}x+9 \right)\left( x^2-3\sqrt{2}x+9 \right)=0\] solve the two quadratics separately by quadratic formula.

OpenStudy (freckles):

cos and sin have period 2pi that is

OpenStudy (freckles):

and you need to take the 4th root of both sides too :)

OpenStudy (freckles):

\[x^4=-81 \\ x^4=81(\cos(\pi)+i \sin(\pi)) \\ \text{ so since \cos and \sin are periodic we have } \\x^4=81(\cos(\pi+2n \pi)+i \sin(\pi+2n \pi)) \] that -81 didn't help much

OpenStudy (anonymous):

geeze. so which strategy do I go with, yours or surj?

OpenStudy (freckles):

either

OpenStudy (anonymous):

ok let me start figuring out

OpenStudy (anonymous):

yay! it worked i went your way, it just seemed easier to me since I am already familiar with it

OpenStudy (freckles):

ok and I changed it from cos(0)+i sin(0) to cos(pi)+isin(pi) because I didn't want to take the square root of -81

OpenStudy (anonymous):

well it worked thank u...

OpenStudy (freckles):

ok

OpenStudy (freckles):

ooops 4th root whatever lol